Two trains left the same city at the same time, traveling in opposite directions. The eastbound train travels for 10 hours, and the westbound train travels for 5 hours. They are now 1300 km apart. The westbound train's rate is 20 km/h faster than the eastbound train. Find the rate speed of each train.
Found 2 solutions by funmath, hotrodd:
Answer by funmath(2933)

Two trains left the same city at the same time, traveling in opposite directions. The eastbound train travels for 10 hours, and the westbound train travels for 5 hours. They are now 1300 km apart. The westbound train's rate is 20 km/h faster than the eastbound train. Find the rate speed of each train.
:
distance = rate*time --> d=rt
The time east t= 10 hr
the rate east r= x
the time west t= 5
the rate west r= x+20
The total distance between them d= 1300 km
:
Therefore it makes sence that if we add the distance that they travel we get 1300 km.
The distance east r*t=10x
the distance west r*t=5(x+20)
:
Probalem to solve:
10x+5(x+20)=1300
10x+5x+100=1300
15x+100-100=1300-100
15x=1200
15x/15=1200/15
x=80
The rate east is x=80 km/h
the rate west is x+20=80+20=100 km/h
Happy Calculating!!!












Answer by hotrodd(18)

The train going westbound is traveling at 100km/h.
The train going eastbound is traveling at 80km/h